INAF queue observing, 2018 Feb 04-05

Observer: F. Cusano, A. Gargiulo
Telescope Operator: D. Gonzalez-Huerta
Support Astronomer: D. Thompson
Instrument:

Summary

At the sunset the sky is clear
instruments: LBC works, LUCI2 is ok, LUCI 1 still MOS problem no AO, MODS2 blue grating only

Night duration (12\xB0 el twilight) 2:00am- 13:15pm UT = 11 h 15m

Observing time: 10h 5m (89.6%)

Weather time loss: 0h m (0%)*

technical time loss 1h 10m (10.4%)

Data Summary

Project_RUN Objects Instrument Filter ExpTime Notes
LBC flat sky flat LBC g,r - B,V   B, V just few are good
LUCI 2 cal dark LUCI2     Prandoni K,J; Saracco
Annibali LBC R17 LBC g, r 0.33h  
Cusano NW-1,2,3 LBC B,V 0.35h up to sequence 18
Rossi GRB171010 LBC r, i 0.5h  
Saracco C1_23152 LUCI2 200g HKspec 2h  
Gilli J LUCI2 200g HKspec 1h completed
Gilli K LUCI2 200g HKspec 2h  
Prandoni USS268 LUCI2 K 15 min  
LUCI2 Sky flat   LUCI2 K,J    

Details

* starting dark with LUCI2
  • dark for Prandoni K luci2.20180205.0001-0005
  • dark for Prandoni J luci2.20180205.0006-0010
  • dark for Saracco luci2.20180205.0011-15 last two darks the dome was opened

01:13 staring sky flat with LBC
  • filter g and r good
  • filter B, V some are good, some have low counts

1:38 slewing to copointing NW-1 Cusano
  • 1:44 dohybrid, problem sending again
  • 1:47 dimm seeing 0.9"
  • 1:49 dohybrid
  • 2:04 red is not converging we stop clear the active optics and run dohybrid again
  • 2:08 dimm seeing 0.8"
  • 2:16 still not converging, we call the SA on call, we clear the active optics and run dofpia
  • 2:26 still problem we move to a different field Annibali R17

2:27 slewing to Annibali R17
  • 2:34 dohybrid
  • 2:45 both sides converged
  • 2:49 lbcrangebal
  • 2:49 sending ob CGCG087_3.ob, R17
  • 2:56 seeing 0.7" on images
  • blue some elongated stars, I have seen already when the seeing is too good, seeing 0.5-0.6" on images
  • images still elongated we will repeat the ob

3:17 slewing to Cusano NW-1
  • 3:20 dofpia
  • 3:26 exec NW-1_18.xml, 3:34 NW-2_18, 3:43 NW-3_18

3:53 slewing to Rossi GRB171010
  • 3:57 dofpia
  • 4:02 copointing ob
  • executing science ob LBT-2017B-C0883_1science.xml
  • 4:11 seeing 0.8" on images
  • 4:20 re-executing science ob LBT-2017B-C0883_1science.xml
  • 4:21 seeing 0.7" on images
  • 4:35 dofpia
  • 4:41 sending again science ob

4:57 switch to LUCI

5:25 slewing to telluric of Saracco
  • 5:25 dimm seeing 0.85"
  • acq luci2.20180205.0016-18
  • 5:57 seeing 0.8" on guider
  • spec 19-22

5:37 slewing to science of Saracco, C1_23152
  • acq 23-26
  • spec 27-50
  • seeing 0.8" on guider
  • seeing 1.1" on guider at 6:21
  • 6:26 seeing 0.9" on guider
  • 6:44 seeing 0.6" on guider
  • 6:57 re-executing science ob Saracco, C1_23152
  • 7:11 seeing 0.5 on guider
  • 7:39 seeing 0.6 on guider

8:05 slewing to telluric of Gilli, Hip 56736
  • acq 51-53
  • spec 54-57

8:19 slewing to Gilli J
  • 8:21 seeing 0.5" on guider
  • acq 58-63
  • spec 64-81
  • 8:41 seeing 0.7" on guider

9:57 slewing to Gilli K
  • acq 82-85
  • spec 86-121
  • 10:10 seeing 0.7" on guider
  • 10:48 seeing 0.7" on guider
  • 11:10 seeing 0.6" on guider
  • 11:24 sending again the spec script, seeing 0.6"
  • 11:48 seeing 0.8" on guider

12:36 slewing to telluric Hip 56736
  • acq 122-124
  • spec 125-128

12:51 slewing to Prandoni USS268 K
  • preset failed, To checks the pointing, 12:57 now is ok
  • first image 129
  • 12:57 seeing 0.6" on guider
  • stopped at 13:21

* starting sky flat LUCI2
  • K band
  • J band
Topic revision: r4 - 07 Feb 2018, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback